OpenStudy (anonymous):

Write the radical expression 8/ 7√x^15 in exponential form.

OpenStudy (anonymous):

do you mean \[\frac{ 8 }{ 7 }\times \sqrt{x ^{15}}\]

OpenStudy (anonymous):

you should know that \[\sqrt[a]{x ^{b}}=x ^{\frac{ b }{ a }}\]

OpenStudy (anonymous):

for example :\[\sqrt[5]{x ^{6}}=x ^{\frac{ 6 }{ 5 }}\]

OpenStudy (triciaal):

the question is not clear?

OpenStudy (triciaal):